OpenStudy (anonymous):

Write an algebraic expression equivalent to cot(arctanx). WILL MEDAL AND FAN

OpenStudy (anonymous):

draw a right triangle

OpenStudy (anonymous):

ok

OpenStudy (anonymous):



OpenStudy (anonymous):

there is a picture of a triangle with an angle labelled, the tangent of that angle is \(x\)

OpenStudy (anonymous):

i was just drawing it lol

OpenStudy (anonymous):

opposite over adjacent

OpenStudy (anonymous):

right

OpenStudy (anonymous):

and the cotangent of that angle should be pretty obvious too

OpenStudy (anonymous):

so im trying to find cotangent

OpenStudy (anonymous):

so 1/x?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

all righty. what about sin(arctanx)?

OpenStudy (anonymous):

same thing

OpenStudy (anonymous):

except this time you need the hypotenuse of the triangle, which you get (in your head) via pythagors

OpenStudy (anonymous):

pythagoras something like that

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then take "opposite over hypotenuse"

OpenStudy (anonymous):

so i get x^2+1^2=c^2

OpenStudy (anonymous):

yikes

OpenStudy (anonymous):

you are solving for the hypotenuse correct?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

i mean you are not wrong, but you should have it instantly one side is 1, one side is x, the hypotenuse is \(\sqrt{1+x^2}\)

OpenStudy (anonymous):

which is what you get when you solve \[x^2+1=c^2\] for \(c\) but you should really just know it

OpenStudy (anonymous):

oh my bad.

OpenStudy (anonymous):

so xsqrt1+x^2/x^2+1??

OpenStudy (anonymous):

my teacher does not like having square roots as the denominator so i put it as the numerator

OpenStudy (anonymous):

ok then pick that one stupid, but ok

OpenStudy (anonymous):

all right so sin(arccos(x+2). i had trouble with this.

OpenStudy (anonymous):

\[x+2^2+B=1\]

OpenStudy (anonymous):

how do i solve that?? @satellite73

OpenStudy (anonymous):

i meant \[x+2^2+B^2=1^2\]

OpenStudy (anonymous):



OpenStudy (anonymous):

\[b=\sqrt{1-(x+2)^2}\] by pythagoras

OpenStudy (anonymous):

you can clean that up a bit if you like or just leave it

OpenStudy (anonymous):

if it were to be cleaned up, would i just simplify the (x+2)^2??

OpenStudy (anonymous):

no

OpenStudy (anonymous):

you would have to expand \((x+2)^2\) and subtract it from \(1\) and leave it inside the radical

OpenStudy (anonymous):

so (x+2)(x+2)-1 all inside the radical?

OpenStudy (anonymous):

i meant 1-(x+2)(x+2)

OpenStudy (anonymous):

yeah or \[\sqrt{1-(4+4x+x^2)}\] or whatever you like

OpenStudy (anonymous):

ok cool. u think you can give me a mini lesson on inverse functions?

OpenStudy (anonymous):

what do you need to know?

OpenStudy (anonymous):

like problems such as arctan(cos1/2). Problems similar to these confuse me.

OpenStudy (anonymous):

most of it is probably right in the book you really need to understand what the range of the inverse functions are (that is what makes them functions) i do those by drawing a triangle,labelling the sides appropriately and then finding the missing side via pythagoras there are other ways, but that is the simplest easiest to answer if you ask a specific question

OpenStudy (anonymous):

EXACTLY. The ranges for the problems confuse me.

OpenStudy (anonymous):

memorize them

OpenStudy (anonymous):

What are the ranges though.

OpenStudy (anonymous):

for arcsine and arctangent it is \[[-\frac{\pi}{2},\frac{\pi}{2}]\]

OpenStudy (anonymous):

I will write that down.

OpenStudy (anonymous):

for arccosine it is \[[0,\pi]\]

OpenStudy (anonymous):

anything else?

OpenStudy (anonymous):

hmm that is the major stuff it will be more complicated when you have to compute something like \[\sin(\tan^{-1}(x)+\cos^{-1}(x))\]

OpenStudy (anonymous):

so what about arcsec or arccosecant or maybe arccotangent. what are the ranges of those?

OpenStudy (anonymous):

yeah im still here lol